OpenStudy (anonymous):

Find the equation of the tangent line to the graph of the given fxn at (0,36) on the graph of the fxn. f(x)=37-sec^3(2x)

OpenStudy (anonymous):

Do you know how to find the derivative of f(x)?

OpenStudy (anonymous):

That's my problem. I think I'm messing up somewhere there.

OpenStudy (anonymous):

\[ f'(x)=-6 \tan (2 x) \sec ^3(2 x) \]

OpenStudy (anonymous):

uhhh

OpenStudy (anonymous):

why isn't it tan^3 and where did you get -6?

OpenStudy (anonymous):

The way to understand it is to put \[ u(x)=\sec(2x)\\ f(x)=37 -u(x)^3\\ f'(x) = - 3 u(x)^2 u'(x) \] Do you understand the steps above?

OpenStudy (anonymous):

now \[ u'(x) =2\sec(2x) \tan(2 x) \] Replace and you are done

OpenStudy (anonymous):

not sure where you got u'(x)

OpenStudy (anonymous):

What is the derivative of sec(2 x)?

OpenStudy (anonymous):

2 cos 2x?

OpenStudy (anonymous):

oh sorry. that was sin 2x

OpenStudy (anonymous):

No. The derivative of sec is sec tan

OpenStudy (anonymous):

ik.

OpenStudy (anonymous):

I meant: f(x)=−3u(x)2u′(x) <---- I thought power rule would just be f(x)=−3u(x)2

OpenStudy (anonymous):

So the derivative of \(\sec(2x)\) is \( 2 \sec(2 x) \tan(2 x)\)

OpenStudy (anonymous):

The power rule is 3 u^2 u'

OpenStudy (anonymous):

How is that possible? http://www.mathwords.com/p/p_assets/p97.gif

OpenStudy (anonymous):

In summary \[ u(x)=\sec(2x)\\ f(x)=37 -u(x)^3\\ f'(x) = - 3 u(x)^2 u'(x)\\ u'(x) =2\sec(2x) \tan(2 x)\\ f'(x)=-6 \tan (2 x) \sec ^3(2 x) \]

OpenStudy (anonymous):

I don't understand that, so could you use the chain rule?

OpenStudy (anonymous):

Remember that the we have here u(x)^3 and not x^3

OpenStudy (anonymous):

Read about the power rule and come back to this problem

OpenStudy (anonymous):

The power rule states that if u is a function of x then the derivative of (\ u^n\) is \( n u^{n-1} u'\)

OpenStudy (anonymous):

Okay. I understand how you did it.

OpenStudy (anonymous):

But can I use the chain rule?

OpenStudy (anonymous):

You are using both the chain rule and the power rule